Martha drew a triangle with side lenghts of 48cm,55cm,and 73cm.Did she draw a right triangle?
Answer by htmentor(1343)

For a right triangle, the Pythagorean theorem says that a^2 + b^2 = c^2
a^2 + b^2 = 48^2 + 55^2 = 5329
c^2 = 73^2 = 5329
So Martha drew a right triangle